# How many measurements are enough? Bayesian recovery in inverse problems with general distributions

### v1

@@ 2.1 Notation
[…]
@@
[…]
@@ 2.2 Approximate covering numbers
As a measure of complexity of measures, we will define the concept of approximate covering numbers as introduced in Jalal et al. jalal2021instance .
@@
  • Remark 2.2 (Relation to non-Bayesian compressed sensing)
    Note that when $\delta=0$, the approximate covering number $\mathrm{Cov}_{\eta,0}(\mathcal{P})\equiv\mathrm{Cov}_{\eta}(\mathrm{supp}(\mathcal{P}))$ is just the classical covering number of the set $\mathrm{supp}(\mathcal{P})$, i.e., the minimal number of balls of radius $\eta$ that cover $\mathrm{supp}(\mathcal{P})$. Classical covering numbers play a key role in (non-Bayesian) compressed sensing theory. Namely, the covering number of the model class $\Sigma\subseteq\mathbb{R}^{n}$ directly determines the number of measurements that suffice for stable and accurate recover. See, e.g., dirksen2016dimensionality ; adcock2024unified . In the Bayesian setting, the approximate covering number plays the same role – see Theorem 1.1.
@@ 2.3 Bounds for $\mathcal{A}$ and $\mathcal{E}$
Let $\mathcal{A}$ be a distribution on $\mathbb{R}^{m\times n}$, $t\geq 0$ and $D\subseteq\mathbb{R}^{n}$. Then a lower concentration bound for $\mathcal{A}$ is any constant $C_{\mathsf{low}}(t)=C_{-}(t;\mathcal{A},D)\geq 0$ such that
@@
$\mathbb{P}({\|Ax\|}>t)\leq C_{\mathsf{abs}}(s,t;\mathcal{A},D),\quad\forall x\in D,\ {\|x\|}\leq s.$
Notice that if $\mathcal{A}$ is isotropic, i.e., $\mathbb{E}{\|Ax\|}^{2}={\|x\|}^{2}$, $\forall x\in\mathbb{R}^{n}$, then $C_{\mathsf{low}}$ and $C_{\mathsf{upp}}$ determine how well ${\|Ax\|}$ concentrates around its mean ${\|x\|}$ for any fixed $x\in D$. In general, one can show that $C_{\mathsf{low}}(t;\mathcal{A},\mathbb{R}^{n}),C_{\mathsf{up}}(t;\mathcal{A},\mathbb{R}^{n})=(t-1)^{-2}=\mathcal{O}(t^{-2})$, $t\rightarrow\infty$. However, in order to obtain desirable sample complexity estimates (e.g., Theorems 1.2 and 1.3), we need to derive concentration bounds that decay exponentially in $m$. A crucial component of this analysis is considering the concentration bound over some subset $D$ (related to the support of $\mathcal{P}$), as, in general, one cannot expect fast concentration over the whole of $\mathbb{R}^{n}$. See Remarks 3 and 3.2.
@@
Notice that this bound just measures the probability that the noise is large. We also need the following bound, which estimates how much the density of $\mathcal{E}$ changes in a $\sigma$-neighbourhood of the origin when perturbed by an amount $\varepsilon$.
@@
Let $\mathcal{E}$ be a distribution in $\mathbb{R}^{m}$ with density $p_{\mathcal{E}}$ and $\varepsilon,\tau\geq 0$. Then a density shift bound for $\mathcal{E}$ is any constant $D_{\mathsf{shift}}(\varepsilon,\tau)=D_{\mathsf{shift}}(\varepsilon,\tau;\mathcal{E})\geq 0$ (possibly +$\infty$+ ∞) such that
@@ 3 Main results
We now present our main results. The first, an extension of Theorem 1.1, is a general result that holds for arbitrary distribution $\mathcal{R}$, $\mathcal{P}$, $\mathcal{A}$ and $\mathcal{E}$.
@@
Let $1\leq p\leq\infty$, $40\leq\delta\leq 1/$ 4, $\varepsilon,\eta,t>0$, $c,c^{\prime}\geq 1$ and $\sigma\geq\varepsilon/\delta^{1/p}$. Let $\mathcal{E}$ be a distribution on $\mathbb{R}^{m}$ and $\mathcal{R},\mathcal{P}$ be distributions in $\mathbb{R}^{n}$ satisfying $W_{p}(\mathcal{R},\mathcal{P})\leq\varepsilon$ and
@@
This theorem bounds the probability of unstable or inaccurate recovery $p$ in terms of the various parameters using the constants introduced in the previous section and the approximate covering numbers of $\mathcal{R}$, $\mathcal{P}$. This result is powerful in its generality, but as a consequence, rather opaque. Later in this section we use it to deduce more explicit results for specific distributions of interest.
  • Remark 3.2 (The concentration bounds in Theorem 3.1)
@@
  • Remark 3.3 (Relation to non-Bayesian compressed sensing)
    The constants $C_{\mathsf{low}}$ and $C_{\mathsf{upp}}$ are similar, albeit not identical to similar conditions such as the Restricted Isometry Property (RIP) (see, e.g., (adcock2021compressive, , Chpt. 5)) or Restricted Eigenvalue Condition (REC) bickel2009simultaneous ; bora2017compressed that appear in non-Bayesian compressed sensing. There, one considers a fixed model class $\Sigma\subseteq\mathbb{R}^{n}$, such as the set $\Sigma_{s}$ of $s$-sparse vectors or, as in bora2017compressed , the range $\mathrm{ran}(G)$ of a generative DNN. Conditions such as the RIP or REC impose that ${\|Ax\|}$ is concentrated above and below around ${\|x\|}$ for all $x$ belonging to the difference set $\Sigma-\Sigma$. In Theorem 3.1, assuming $\mathcal{P}$ attains the minimum in (3.1), there is a similar condition with $\Sigma$ replaced by $\mathrm{supp}(\mathcal{P})$. Indeed, $C_{\mathsf{low}}(2\sqrt{2}/\sqrt{2};\mathcal{A},D_{2})$ measures how small ${\|Ax\|}$ is in relation to ${\|x\|}$, with $C_{\mathsf{upp}}(\sqrt{c}/(2\sqrt{2});\mathcal{A},D_{2})$ measuring how large ${\|Ax\|}$ is in relation to ${\|x\|}$.
@@ 3.1 Example: Subgaussian random matrices with Gaussian noise
Note that $1/\sqrt{m}$ is a scaling factor which ensures that $A$ is isotropic, i.e., $\mathbb{E}{\|Ax\|}^{2}={\|x\|}^{2}$, $\forall x\in\mathbb{R}^{n}$.
@@
Let $1\leq p\leq\infty$, $40\leq\delta\leq 1/$ 4, $\varepsilon,\eta>0$ and $\sigma\geq\varepsilon/\delta^{1/p}$. Let $\mathcal{E}=\mathcal{N}(0,\frac{\sigma^{2}}{m}I)$ and $\mathcal{A}$ be a distribution of subgaussian random matrices with parameters $\beta,\kappa>0$. Suppose that $x^{*}\sim\mathcal{R}$, $A\sim\mathcal{A}$, $e\sim\mathcal{E}$ independently and $\hat{x}\sim\mathcal{P}(\cdot|y,A)$, where $y=Ax^{*}+e$. Then there is a constant $c(\beta,\kappa)>0$ depending on $\beta,\kappa$ only such that
@@
provided
@@ 3.2 Example: randomly-subsampled orthogonal transforms with Gaussian noise
As discussed, subgaussian random matrices are largely impractical. We now discuss the more practical case of subsampled orthogonal transforms.
@@
Let $U\in\mathbb{R}^{n\times n}$ be orthogonal (i.e., $U^{\top}U=UU^{\top}=I$). and write $u_{1},\ldots,u_{n}\in\mathbb{R}^{n}$ for its rows. Let $X_{1},\ldots,X_{n}\sim_{\mathrm{i.i.d.}}\mathrm{Ber}(m/n)$ be independent Bernoulli random variables with $\mathbb{P}(X_{i}=1)=m/n$ and $\mathbb{P}(X_{i}=0)=1-m/n$. Then we define a distribution $\mathcal{A}$ as follows. We say that $A\sim\mathcal{A}$ if
@@
The factor $\sqrt{n/m}$ ensures that $\mathbb{E}(A^{*}A)=I$. Note that the number of measurements $q$ in this model is itself a random variable, with $\mathbb{E}(q)=m$. However, $q$ concentrates exponentially around its mean.
@@
Coherence is a well-known concept in classical compressed sensing with sparse vectors. Definition 3.7 is a generalization that allows for arbitrary model classes $\Sigma$. This definition is similar to that of berk2022coherence , which considered non-Bayesian compressed sensing with generative models. It is related to the (somewhat more general) concept of variation introduced in adcock2024unified .
@@
Let $1\leq p\leq\infty$, $40\leq\delta\leq 1/$ 4, $\varepsilon,\eta>0$ and $\sigma\geq\varepsilon/\delta^{1/p}$. Let $\mathcal{E}=\mathcal{N}(0,\frac{\sigma^{2}}{m}I)$ and $\mathcal{A}$ be a distribution of randomly-subsampled orthogonal matrices based on a matrix $U$. Suppose that $x^{*}\sim\mathcal{R}$, $A\sim\mathcal{A}$, $e\sim\mathcal{E}$ independently and $\hat{x}\sim\mathcal{P}(\cdot|y,A)$, where $y=Ax^{*}+e$. Then there is a universal constant $c>0$ such that
@@
provided
@@
This theorem can be considered a generalization of results from berk2022coherence ; adcock2024unified to the Bayesian setting. In berk2022coherence ; adcock2024unified , the measurement conditions scale linearly with $\mu(U;D)$, where $D=\Sigma-\Sigma$ and $\Sigma\subseteq\mathbb{R}^{n}$ is the low-complexity model class. Similarly, the number of measurements (3.5) scales linearly with respect to the coherence relative to $D=\mathrm{supp}(\mathcal{P})-\mathrm{supp}(\mathcal{P})$, which, as discussed in Remark 3, plays the role of the low-complexity model class in the Bayesian setting.
It is known in classical compressed sensing that coherence determines the sample complexity of recovering sparse vectors from randomly-subsampled orthogonal transforms candes2011probabilistic . A similar argument can be made in the Bayesian setting. Notice that $\mu(U;D)\leq\mu(U;\mathbb{R}^{n})=n$. However, we are particularly interested in cases where $\mu(U;D)\ll n$, in which case (3.5) may be significantly smaller than the ambient dimension $n$. We discuss this in the context of several examples in the next section.
  • Remark 3.9 (Concentration over subsets)

### v2

@@ 2.1 Notation
We now introduce some further notation. We let $B_{r}(x)=\{z\in\mathbb{R}^{n}:{\|z-x\|}\leq r\}$ and, when $x=0$, we write $B_{r}:=B_{r}(0)$. Given a set $X\subseteq\mathbb{R}^{n}$, we write $\XX^{c}=\mathbb{R}^{n}\backslash X$ for its complement. We also write $B_{r}(X)=\bigcup_{x\in X}B_{r}(x)$ for the $r$-neighbourhood of $X$.
@@
$W_{p}(\mu,\nu)=\inf_{\gamma\in\Gamma}\left(\int_{X\times Y}c(x,y)^{p}d\gamma(x,y)\right)^{1/p}.$
@@ 2.2 Approximate covering numbers
As a measure of complexity of measures, we use the concept of approximate covering numbers as introduced in jalal2021instance .
@@
Note that when $\delta=0$, the approximate covering number $\mathrm{Cov}_{\eta,0}(\mathcal{P})\equiv\mathrm{Cov}_{\eta}(\mathrm{supp}(\mathcal{P}))$ is just the classical covering number of the set $\mathrm{supp}(\mathcal{P})$, i.e., the minimal number of balls of radius $\eta$ that cover $\mathrm{supp}(\mathcal{P})$. Classical covering numbers play a key role in (non-Bayesian) compressed sensing theory. Namely, the covering number of (the unit ball of) the model class $\Sigma\subseteq\mathbb{R}^{n}$ directly determines the number of measurements that suffice for stable and accurate recovery. See, e.g., dirksen2016dimensionality ; adcock2024unified . In the Bayesian setting, the approximate covering number plays the same role; see Theorem 1.1.
@@ 2.3 Bounds for $\mathcal{A}$ and $\mathcal{E}$
Let $\mathcal{A}$ be a distribution on $\mathbb{R}^{m\times n}$, $t\geq 0$ and $D\subseteq\mathbb{R}^{n}$. Then a lower concentration bound for $\mathcal{A}$ is any constant $C_{\mathsf{low}}(t)=C_{\mathsf{low}}(t;\mathcal{A},D)\geq 0$ such that
@@
$\mathbb{P}_{A\sim\mathcal{A}}({\|Ax\|}>t)\leq C_{\mathsf{abs}}(s,t;\mathcal{A},D),\quad\forall x\in D,\ {\|x\|}\leq s.$
Notice that if $\mathcal{A}$ is isotropic, i.e., $\mathbb{E}{\|Ax\|}^{2}={\|x\|}^{2}$, $\forall x\in\mathbb{R}^{n}$, then $C_{\mathsf{low}}$ and $C_{\mathsf{upp}}$ determine how well ${\|Ax\|}^{2}$ concentrates around its mean ${\|x\|}^{2}$ for any fixed $x\in D$. To obtain desirable sample complexity estimates (e.g., Theorems 1.2 and 1.3), we need concentration bounds that decay exponentially in $m$. A crucial component of this analysis is considering concentration bounds over some subset $D$ (related to the support of $\mathcal{P}$), as, in general, one cannot expect fast concentration over the whole of $\mathbb{R}^{n}$. See Remarks 3 and 3.2.
@@
Notice that this bound just measures the probability that the noise is large. We also need the following concept, which estimates how much the density of $\mathcal{E}$ changes in a $\tau$-neighbourhood of the origin when perturbed by an amount $\varepsilon$.
@@
Let $\mathcal{E}$ be a distribution in $\mathbb{R}^{m}$ with density $p_{\mathcal{E}}$ and $\varepsilon,\tau\geq 0$. Then a density shift bound for $\mathcal{E}$ is any constant $D_{\mathsf{shift}}(\varepsilon,\tau)=D_{\mathsf{shift}}(\varepsilon,\tau;\mathcal{E})\geq 0$ (possibly $\infty$) such that
@@ 3 Main results
We now present our main results. The first, an extension of Theorem 1.1, is a general result that holds for arbitrary distributions $\mathcal{R}$, $\mathcal{P}$, $\mathcal{A}$ and $\mathcal{E}$.
@@
Let $1\leq p\leq\infty$, $0\leq\delta\leq 1/4$, $\varepsilon,\eta,t>0$, $c,c^{\prime}\geq 1$ and $\sigma\geq\varepsilon/\delta^{1/p}$. Let $\mathcal{E}$ be a distribution on $\mathbb{R}^{m}$ and $\mathcal{R},\mathcal{P}$ be distributions on $\mathbb{R}^{n}$ satisfying $W_{p}(\mathcal{R},\mathcal{P})\leq\varepsilon$ and
@@
This theorem bounds the probability $p$ of unstable or inaccurate recovery in terms of the various parameters using the constants introduced in the previous section and the approximate covering numbers of $\mathcal{R}$, $\mathcal{P}$. This result is powerful in its generality, but as a consequence, rather opaque. In particular, since it considers arbitrary measurement and noise distributions, the number of measurements $m$ does not explicitly enter the bound. For typical distributions, a dependence on $m$ is found in the concentration bounds $C_{\mathsf{low}},C_{\mathsf{upp}},D_{\mathsf{upp}}$, as well as the terms $D_{\mathsf{shift}}$ and $C_{\mathsf{abs}}$. For instance, the former decay exponentially-fast in $m$ for the examples introduced in §1.2, and therefore compensate for the exponentially-large scaling in $k$ in the main bound (see §B for precise estimates, as well as the discussion in §3.1-3.2). Note that Theorem 3.1 also considers general noise distributions $\mathcal{E}$. While Gaussian noise is arguably the most important example – and will be used in all our subsequent examples – this additional generality comes at little cost in terms of the technicality of the proofs.
@@
The constants $C_{\mathsf{low}}$ and $C_{\mathsf{upp}}$ are similar, albeit not identical to similar conditions such as the Restricted Isometry Property (RIP) (see, e.g., (adcock2021compressive, , Chpt. 5)) or Restricted Eigenvalue Condition (REC) bickel2009simultaneous ; bora2017compressed that appear in non-Bayesian compressed sensing. There, one considers a fixed model class $\Sigma\subseteq\mathbb{R}^{n}$, such as the set $\Sigma_{s}$ of $s$-sparse vectors or, as in bora2017compressed , the range $\mathrm{ran}(G)$ of a generative DNN. Conditions such as the RIP or REC impose that $\|Ax\|$} is concentrated around $\|x\|$} for all $x$ belonging to the difference set $\Sigma-\Sigma$. In Theorem 3.1, assuming $\mathcal{P}$ attains the minimum in (3.1), there is a similar condition with $\Sigma$ replaced by $\mathrm{supp}(\mathcal{P})$. Indeed, $C_{\mathsf{low}}(2\sqrt{2}/\sqrt{c};\mathcal{A},D_{2})$ measures how small $\|Ax\|$} is in relation to $\|x\|$} and $C_{\mathsf{upp}}(\sqrt{c}/(2\sqrt{2});\mathcal{A},D_{2})$ measures how large $\|Ax\|$} is in relation to $\|x\|$}.
@@ 3.1 Example: Subgaussian random matrices with Gaussian noise
Note that $1/\sqrt{m}$ is a scaling factor that ensures that $A$ is isotropic, i.e., $\mathbb{E}{\|Ax\|}^{2}={\|x\|}^{2}$, $\forall x\in\mathbb{R}^{n}$.
@@
Let $1\leq p\leq\infty$, $0\leq\delta\leq 1/4$, $\varepsilon,\eta>0$ and $\sigma\geq\varepsilon/\delta^{1/p}$. Let $\mathcal{E}=\mathcal{N}(0,\frac{\sigma^{2}}{m}I)$ and $\mathcal{A}$ be a distribution of subgaussian random matrices with parameters $\beta,\kappa>0$. Let $\mathcal{R},\mathcal{P}$ be distributions on $\mathbb{R}^{n}$ and suppose that $x^{*}\sim\mathcal{R}$, $A\sim\mathcal{A}$, $e\sim\mathcal{E}$ independently and $\hat{x}\sim\mathcal{P}(\cdot|y,A)$, where $y=Ax^{*}+e$. Then there is a constant $c(\beta,\kappa)>0$ depending on $\beta,\kappa$ only such that
@@
provided $W_{p}(\mathcal{R},\mathcal{P})\leq\varepsilon/c(\beta,\kappa)$ and
@@
It is worth at this stage discussing how (3.4) behaves with respect to the various parameters. First, suppose that $\eta$ decreases so that the error bound becomes smaller. Then $\mathrm{Cov}_{\eta,\delta}(\cdot)$ increases, meaning, as expected, that more measurements are required to meet (3.4). Second, suppose that $\delta$ decreases, so that the failure probability shrinks. Then $\mathrm{Cov}_{\eta,\delta}(\cdot)$ and $\log(1/\delta)$ both increase, meaning, once again, that more measurements are needed for (3.4) to hold. Both behaviours are as expected.
Suppose that $\mathcal{P}=\mathcal{R}$ is a sum of $d$ Diracs located at $X=\{x_{1},\ldots,x_{d}\}\subseteq\mathbb{R}^{n}$. Since the matrix $A\in\mathbb{R}^{m\times n}$ is a linear dimensionality-reducing map, the Johnson-Lindenstrauss Lemma states that distances in $X$ are preserved under $A$ if and only if $m=O(\log(d))$. In this setting, preserving the distances in $X$ is equivalent to being able to stably identify the mode from which a signal is drawn when observing its measurements. In agreement with this argument, Theorem 3.5 also predicts recovery from roughly $m=O(\log(d))$ measurements.
@@ 3.2 Example: Randomly-subsampled orthogonal transforms with Gaussian noise
As discussed, subgaussian random matrices are largely impractical. We now consider the more practical case of subsampled orthogonal transforms.
@@
Let $U\in\mathbb{R}^{n\times n}$ be orthogonal (i.e., $U^{\top}U=UU^{\top}=I$) and write $u_{1},\ldots,u_{n}\in\mathbb{R}^{n}$ for its rows. Let $X_{1},\ldots,X_{n}\sim_{\mathrm{i.i.d.}}\mathrm{Ber}(m/n)$ be independent Bernoulli random variables with $\mathbb{P}(X_{i}=1)=m/n$ and $\mathbb{P}(X_{i}=0)=1-m/n$. Then we define a distribution $\mathcal{A}$ as follows. We say that $A\sim\mathcal{A}$ if
@@
The factor $\sqrt{n/m}$ ensures that $\mathbb{E}(A^{\top}A)=I$. Note that the number of measurements $q$ in this model is itself a random variable, with $\mathbb{E}(q)=m$. However, $q$ concentrates exponentially around its mean.
@@
Coherence is a well-known concept in classical compressed sensing with sparse vectors. Definition 3.8 is a generalization that allows for arbitrary model classes $D$. This definition is similar to that of berk2022coherence , which considered non-Bayesian compressed sensing with generative models. It is also related to the more general concept of variation introduced in adcock2024unified .
@@
Let $1\leq p\leq\infty$, $0\leq\delta\leq 1/4$, $\varepsilon,\eta>0$ and $\sigma\geq\varepsilon/\delta^{1/p}$. Let $\mathcal{E}=\mathcal{N}(0,\frac{\sigma^{2}}{m}I)$ and $\mathcal{A}$ be a distribution of randomly-subsampled orthogonal matrices based on a matrix $U$. Let $\mathcal{R},\mathcal{P}$ be distributions on $\mathbb{R}^{n}$ and suppose that $x^{*}\sim\mathcal{R}$, $A\sim\mathcal{A}$, $e\sim\mathcal{E}$ independently and $\hat{x}\sim\mathcal{P}(\cdot|y,A)$, where $y=Ax^{*}+e$. Then there is a universal constant $c>0$ such that
@@
provided $W_{p}(\mathcal{R},\mathcal{P})\leq\varepsilon/(2\sqrt{mn})$ and
@@
This theorem is a Bayesian analogue of the deterministic results shown in berk2022coherence ; adcock2024unified . In berk2022coherence ; adcock2024unified , the measurement conditions scale linearly with $\mu(U;D)$, where $D=\Sigma-\Sigma$ and $\Sigma\subseteq\mathbb{R}^{n}$ is the low-complexity model class. Similarly, the number of measurements (3.5) scales linearly with respect to the coherence relative to $D=\mathrm{supp}(\mathcal{P})-\mathrm{supp}(\mathcal{P})$, which, as discussed in Remark 3, plays the role of the low-complexity model class in the Bayesian setting. Note that the measurement condition (3.5) involves the approximate distribution $\mathcal{P}$ only. This is relevant, since the quantities $\mu(U;D)$ and $\mathrm{Cov}_{\eta,\delta}(\mathcal{P})$ can be estimated either numerically or analytically in various cases, such as when $\mathcal{P}$ is given by a generative model. Indeed, we estimate $\mathrm{Cov}_{\eta,\delta}(\mathcal{P})$ analytically for Lipschitz generative models in Proposition 4.1 below. The coherence $\mu(U;D)$ was estimated analytically in berk2022coherence for ReLU DNNs with random weights (see Remark 4.1 below). It can also be estimated numerically for more general types of generative models berk2022coherence ; adcock2023cs4ml . Overall, by estimating these quantities, one can use (3.5) to gauge how well one can recover with a given $\mathcal{P}$. Note that this may not be possible if (3.5) involved $\mathcal{R}$ as well, since this distribution is typically unknown.
In classical compressed sensing, coherence determines the sample complexity of recovering sparse vectors from randomly-subsampled orthogonal transforms candes2011probabilistic . A similar argument can be made in the Bayesian setting. Notice that $\mu(U;D)\leq\mu(U;\mathbb{R}^{n})=n$. However, we are particularly interested in cases where $\mu(U;D)\ll n$, in which case (3.5) may be significantly smaller than the ambient dimension $n$. We discuss this in the context of several examples in the next section.